OpenStudy (anonymous):

0 < x < 2pi , cot theta= squareroot 3

OpenStudy (anonymous):

\[\cot \theta = \sqrt{3}\]

OpenStudy (ash2326):

\(\tan \) and \(\cot\) are positive in the first and third quadrant. That's \[0\le \theta \le \pi/2\ and\ \pi\le \theta \le 3\pi/2\] Do you get this?

OpenStudy (ash2326):

@burhan101 ?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

For what value of theta cot \(\theta\) = \(\sqrt 3\) ?

OpenStudy (ash2326):

tan \(\theta\) = \(\frac 1{\sqrt 3}\) for \(\theta\)= 30 degrees. Do you know when what's \(\cot 30\)?

OpenStudy (anonymous):

-tan

OpenStudy (ash2326):

oops, what's the relation between tan and cot?

OpenStudy (anonymous):

inverse of each other

OpenStudy (ash2326):

so if \[\tan 30= \sqrt 3\] then what's \[\cot 30=??\]

OpenStudy (anonymous):

-\[\sqrt{3}\]

OpenStudy (ash2326):

sorry \[\tan 30=\frac 1{\sqrt 3}\] then what's \(\cot 30\)???

OpenStudy (anonymous):

\[\sqrt{3}\]

OpenStudy (ash2326):

good so we found one answer so \[30\ degrees= \frac \pi {6}\]

OpenStudy (ash2326):

ok tell me what's? \[\cot (180+x)\ or\ \cot (\pi+x)\]

OpenStudy (anonymous):

30

OpenStudy (ash2326):

nope, it won't be that. we know that tan/ cot are positive in first and third quadrant so \[\tan (\pi+x)=\tan x\] so what's \[\cot (\pi+x)=??\]

OpenStudy (ash2326):

@burhan101 ??